OpenStudy (anonymous):

Please help Uptown Photography charges $15 per portrait sheet, plus $19 for a session fee. Downtown Photography charges $29 for a session fee and $10 per portrait sheet. Find the number of portrait sheets for which the total price is equal for the companies with the solution stated as an ordered pair.

imqwerty (imqwerty):

let the number of portraits be n total charge by uptown photography=15n+19 total charge by downtown photography=10n+29 total charge by uptown photography=total charge by downtown photography 15n+19=10n+29 15n-10n=29-19 5n=10 n=2 so 2 photos

OpenStudy (anonymous):

Thank you.

imqwerty (imqwerty):

np :)